U(x1​,x2​)="(\sqrt{x_1}+\sqrt{x_2})^2"


1
Expert's answer
2021-09-20T16:37:23-0400

Find the partial derivatives of X1 and X2 using the utility function

U("X_{1}" ​,"X_{2}" ​)= "( x_1^{0.5} + x_2^{0.5} )^{2}"

Using chain rule we partially differentiate the utility function with respect to "X_{1} and X_{2}"

Let Z = "x_1^{0.5} + x_2^{0.5}"


"\frac{ɗz}{ɗx_{1}} = \frac{0.5}{X_{1}^{0.5}}"


"\frac{ɗz}{ɗx_{2}} = \frac{0.5}{X_{2}^{0.5}}"


"U(X_{1}\n\u200b\n \u200b,X_{2}\n\u200b\n \u200b)" = Z"^{2}"


"\frac{ɗu}{ɗz} =" 2Z


"\frac{ɗu}{ɗx_{1}} = \frac{0.5}{X_{1}^{0.5}} \times 2[x_1^{0.5} + x_2^{0.5}]"


"\frac{ɗu}{ɗx_{1}} = 1 + 2x_2^{0.5}"


"\frac{ɗu}{ɗx_{2}} = \frac{0.5}{X_{2}^{0.5}} \times 2[x_1^{0.5} + x_2^{0.5}]"


"\frac{ɗu}{ɗx_{2}} = 1 + 2x_1^{0.5}"